You form a portfolio of stocks. Stock A has an expected return of 2.8%, Stock B has an expected return of 4.4%, and Stock C has an expected return of 7.6%. If 6% of your portfolio is invested in stock A, and 35% of your stock is invested in Stock C, what is your expected portfolio return?

(Enter your response as a percentage with two decimal places, ex: 12.34)
Answer #1

Investment in B=100-(6+35)=59%

Portfolio return=Respective return*Respective weight

=(2.8*0.06)+(4.4*0.59)+(7.6*0.35)

which is equal to

=5.42%(Approx).